Science Operations

Recent Announcements

WCS added to MODS headers

Posted Jul 08, 2022

To satisfy the requirements that LBTO data archived and distributed to partners be IVOA-compliant, a WCS was
added to MODS headers at the end of April 2022. MODS data taken on and after 2022-04-29 UT will contain a
spatial WCS …

>> Read More

PEPSI Use May9 – June 30

Posted May 12, 2022

Regarding PEPSI scheduling for the remainder of 2022A, the PEPSI team will be conducting their fourth and final run using the 1.8 m VATT plus fiber feed to PEPSI for their TESS campaign tied to ecliptic pole visibility. Their …

>> Read More

LUCI1 MOS failure – Resolved

Posted Mar 28, 2022

The LUCI1 1.5/2.0 Longslit mask experienced a MOS mask failure moving from turnout to FPU.  During this failure the mask rotator failed to move and the mask stuck in an intermediate state.  This MOS error could be caused by loose …

>> Read More

MODS2 Cassette Mechanism Damage – Partially Resolved

Posted Mar 01, 2022

During operations last night there was a MODS slitmask fault error:
2022-03-01T06:08:37.819582 M2.IE>MC2 ERROR: SLITMASK SLITMASK=FAULT Mechanism out-of-position, in-position sensor not asserted. Reset SLITMASK to recover
Visual inspection of the mask loading port showed damage to the upper part of …

>> Read More

LUCI1 AO Patrol Field Change

Posted Feb 21, 2022

The LUCI1 AO Patrol field is pending remeasurement after a recent intervention. The field has slightly changed. Users should avoid selecting stars too close to the edge of the field as they appear in the OT as there …

>> Read More

MODS2 Ne Lamp Failing to fire – Resolved

Posted Feb 17, 2022

The MODS Neon Lamp is failing to fire consistently. It appears to be the result of a bad power supply. A new power supply has been obtained and is awaiting installation. (Update on 3-Mar-2022: The power supplies and Ne lamps …

>> Read More